OpenStudy (anonymous):

when solving for x in ln(x+5)+ln(x^2)=ln(6x) can one just divide both sides by ln to obtain x^2-5x+5?

OpenStudy (anonymous):

Solve with some x and see......

OpenStudy (anonymous):

The roots of the quadratic are 1/2 (5 plus/minus sqrt5) so if you calculate with those in the ln expression, you will see the answer to your question.

OpenStudy (calculusfunctions):

I'm afraid @Razzputin and @estudier, those answers are not correct.

OpenStudy (calculusfunctions):

Would you like me to help?

OpenStudy (anonymous):

I agree that I could have just said no, if that's what u mean....:-)

OpenStudy (calculusfunctions):

@estudier, no, all I meant was that the answers I see are not right. I'm not sure what you're talking about?

OpenStudy (anonymous):

The question is " can one just divide both sides by ln " ?

OpenStudy (anonymous):

Poster did not ask for a solution.

OpenStudy (calculusfunctions):

On the left side, you should first apply the following property:\[\ln xy =\ln x +\ln y\]where both x and y are greater than zero.

OpenStudy (calculusfunctions):

I know he didn't ask for a solution and that's clearly not what I'm doing. But he obviously needs some kind of lesson to achieve some level of understanding because he wrongly believes that you can divide by ln which is NOT true. Now I think that you are implying that my help is not needed here so that's fine. I'll leave. Sorry to have bothered you @estudier.

OpenStudy (calculusfunctions):

If @Razzputin asks for my help then I'll be more than happy to. Simply let me know. Good Luck!

OpenStudy (anonymous):

I'm not implying anything, I am just explaining to you that my answer is not "incorrect" as you stated..-)

OpenStudy (calculusfunctions):

Would you like to see my explanation and why I believe that your answer is incorrect?

OpenStudy (anonymous):

No, thank you.

OpenStudy (anonymous):

sorry i disapeared im back now had to have lunch with my dad

OpenStudy (calculusfunctions):

@Razzputin, do you see the property I gave above?

OpenStudy (anonymous):

yes i do. Once i use that then do i divide both sides?

OpenStudy (anonymous):

ln(x^3+5x^2)=ln(6x) x^3+5x^2-6x=0

OpenStudy (calculusfunctions):

NO! Can you please show me what you have after you apply the property so that I can guide you further? I just don't want to end up giving you the answer. I want to teach you, if you could please show me your first step. Oh I see you did! That's correct, now what do you think you should do next? Can you tell me?

OpenStudy (calculusfunctions):

OK, so what should you do next?

OpenStudy (anonymous):

\[x(x ^{2}+5x-6)=0\] x(x-1)(x+6)=0 x=0 x=1 x=-6 only x=1 is applicable as in logs x>0

OpenStudy (calculusfunctions):

Yes and now what?

OpenStudy (anonymous):

lol NO! .... YES! confusion in the ranks ha ha.

OpenStudy (calculusfunctions):

Perfect! Only x = 1 is a valid solution. Do you now have a better understanding of how to solve these types of problems?

OpenStudy (anonymous):

yup yhanks, i see now that i can only divide the two sides if only one ln is present on each after simplifying them

OpenStudy (calculusfunctions):

@Razzputin and @hartnn, thank you for the medal!

OpenStudy (anonymous):

no prob u deserved it. Most ppl here just hand out the answer its anoying

hartnn (hartnn):

just a side note when u got x^3+5x^2-6x=0 , did u say, u divide both sides ?? what by ln ?? *one cannot simply divide by ln on both side*, its like you are equating the argument of ln lnA = ln B then A+B

hartnn (hartnn):

A=B

OpenStudy (calculusfunctions):

If \[\ln f(x)=\ln g(x)\]then f(x) = g(x) as long as f(x) and g(x) are both greater than zero.

OpenStudy (anonymous):

well i had ln(x^3+5x^2 )=ln(6x) so i dropped the ln

hartnn (hartnn):

yeah, but u DID NOT DIVIDE by ln.....keep that in mind...

OpenStudy (calculusfunctions):

NO! dividing by ln is not allowed.

OpenStudy (anonymous):

lol my bad, ha ha wrong terminology

OpenStudy (calculusfunctions):

If you have two logarithms to the same base that are equal, then you equate their arguments (antilogarithms), but it is wrong to say that you divide by ln. Understand?

OpenStudy (anonymous):

yup thanks

OpenStudy (calculusfunctions):

Welcome! No need to even say thanks. I love teaching and I love math!